THe problem states: Solve by completing the square x^2 - 8x + 4 = 10.
Answer by Nate(3500)

x^2 - 8x + 4 = 10
x^2 - 8x = 6
(x - 4)^2 = 22
x - 4 = +- sqrt(22)
x = 4 +- sqrt(22)